OpenStudy (anonymous):

Find x: 2x^2- 7x -9=0

OpenStudy (raden):

use factorization : 2x^2- 7x -9=0 (2x - 9)(x + 1) = 0 solve for x by take the zeroes of all term be equal 0 2x-9=0 ----> x= ? x+1=0 -----> x= ?

OpenStudy (anonymous):

another way to do it is using the quadratic formula

OpenStudy (anonymous):

Can you solve it for me using the quadratic formula so I can see